OpenStudy (anonymous):

The variables x and y vary directly. Given x = 5 and y = 3, write an equation that relates the variables and find y when x = 2. Answer y = 15x; y = 30 y = 5/3 x; y = 10/3 y= 0.6x; x = 1.2 y = 15/x; y = 7.5

OpenStudy (callisto):

Since variables x and y vary directly, \[\frac{x_{1}}{y_{1}} = \frac{x_2}{y_2}\]\[\frac{5}{3} = \frac{2}{y}\] Can you continue?

OpenStudy (anonymous):

no is it cross multiply

OpenStudy (callisto):

Yes. What would you get then?

OpenStudy (anonymous):

its B

OpenStudy (callisto):

No... Can you show your workings?

OpenStudy (anonymous):

i just got 10/3

OpenStudy (callisto):

Steps?

OpenStudy (anonymous):

idk im stumped

OpenStudy (callisto):